OpenStudy (love_to_love_you):

Okay Idk if I did this one right @loser66 f(x) = (5x - 15) / 4

OpenStudy (love_to_love_you):

f^-1(x)= 3 + 4x/5

OpenStudy (love_to_love_you):



zepdrix (zepdrix):

What are you trying to do? Find the inverse function from this? \(\large\rm f(x)=\frac{5x-15}{4}\) and then graph it?

satellite73 (satellite73):

is the question to find the inverse?

satellite73 (satellite73):

if so, your answer is correct \[f^{-1}(x)=3+\frac{4x}{5}\]

OpenStudy (loser66):

perfect!!